$100 is deposited at the end of every week for five years in an account that pays 14%/a, compounded weekly. 

a. What type of annuity is this?

b. Find the future value of the annuity using the formula.

c. Find the future value of the annuity using a spreadsheet. In your answer include the formulas that you typed.

d. Find the future value of the annuity using the TVM solver on a graphing calculator or on a website. In your answer, include what values you typed for each parameter. 
Answer #1

a) Ordinary weekly annuity

b)

PMT = 100

r = 14%/52 = 0.0026923 (since there are 52 weeks in an year)

n = 5*52 = 260 weeks

(1+r) - 1 FV = PMT * T

FV = 100*[((1+0.0026923)^260 )-1)/0.0026923 ]

FV = 37583.3

c)

A B C 1 0.00269231 2 3 Rate nper PMT PV FV 260 100 0 $37,583.35 4 5 6

A 1 2. 3 Rate nper PMT B =0.14/52 1=5*52 =100 =0 =-FV(B1,B2,B3,B4) 4 PV 5 FV LD 6

d)

Using a financial calculator

PV = 0

PMT = 100

N = 5*52

I/Y = 14/52

cpt FV, we get FV = 37583.3